State the Pigeonhole Principle. Prove that if six integers are selected from the set

[3,4,5,6,7,8,9,10,11,12] there must be two integer whose sum is fifteen.


1
Expert's answer
2021-11-23T18:03:17-0500

For natural numbers "{\displaystyle k}"  and "{\displaystyle m}," if "{\displaystyle n=km+1}" objects are distributed among "{\displaystyle m}" sets, then the Pigeonhole Principle asserts that at least one of the sets will contain at least "{\displaystyle k+1}"  objects.


Let us consider the partition of the set "\{3,4,5,6,7,8,9,10,11,12\}" ito 5 subset: "\{3,12\}""\{4,11\}""\{5,10\}," "\{6,9\}" and "\{7,8\}." The sum of elements in each subset is equal to 15. Taking into account that there are 5 subsets and we choose 6 elements by Pigeonhole Principle we get that at least one pair will be selected. Consequently, there must be two integer whose sum is fifteen.